# Percolation on graphs of polynomial growth is local: analyticity, supercritical sharpness, isoperimetry

### v1

@@ 6 Investigating $\theta$ near $p_{c}$
The first condition implies the second one because Proposition 1.9 asserts that its conditions 1 and 5 are equivalent. We now prove the reverse implication. Let $\mathscr{G}\in\mathfrak{A}$. Construct $\mathscr{G}_{n}$ as in the proof of Proposition 1.9. Observe that $\mathscr{G}_{n}$ is a quotient of $\mathscr{G}$, so that Proposition 1 from [BS96] guarantees that $p_{c}(\mathscr{G}_{n})\geqslant p_{c}(\mathscr{G}_{\infty})$. This means that the domain of $\tilde{\theta}_{\mathscr{G}_{\infty}}$ contains that of $\tilde{\theta}_{\mathscr{G}_{n}}$. As [DCST16] guarantees that $\theta_{\mathscr{G}_{n}}$ is continuous, we have $\inf_{p>p_{c}(\mathscr{G}_{n})}\tilde{\theta}_{\mathscr{G}_{n}}(x)=0$. If we assume uniform convergence of $\tilde{\theta}_{\mathscr{G}_{n}}$ to $\tilde{\theta}_{\mathscr{G}_{\infty}}$, this yields $\inf_{p>p_{c}(\mathscr{G})}\tilde{\theta}_{\mathscr{G}}(x)\leqslant 0$, hence that $\mathscr{G}$ satisfies $\theta(p_{c})=0$. ∎

### v2

@@ 6 Investigating $\theta$ near $p_{c}$
The first condition implies the second one because Proposition 1.10 asserts that its conditions 1 and 5 are equivalent. We now prove the reverse implication. Let $\mathscr{G}\in\mathfrak{A}$. Construct $\mathscr{G}_{n}$ as in the proof of Proposition 1.10. Observe that $\mathscr{G}_{n}$ is a quotient of $\mathscr{G}$, so that Proposition 1 from [BS96] guarantees that $p_{c}(\mathscr{G}_{n})\geqslant p_{c}(\mathscr{G}_{\infty})$. This means that the domain of $\tilde{\theta}_{\mathscr{G}_{\infty}}$ contains that of $\tilde{\theta}_{\mathscr{G}_{n}}$. As [DST16] guarantees that $\theta_{\mathscr{G}_{n}}$ is continuous, we have $\inf_{p>p_{c}(\mathscr{G}_{n})}\tilde{\theta}_{\mathscr{G}_{n}}(x)=0$. If we assume uniform convergence of $\tilde{\theta}_{\mathscr{G}_{n}}$ to $\tilde{\theta}_{\mathscr{G}_{\infty}}$, this yields $\inf_{p>p_{c}(\mathscr{G})}\tilde{\theta}_{\mathscr{G}}(x)\leqslant 0$, hence that $\mathscr{G}$ satisfies $\theta(p_{c})=0$. ∎